Circular No. 5640

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992be IN ANONYMOUS GALAXY Jean Mueller reports her discovery of a supernova of mag ∼ 17 on a

red plate taken Sept. 30 and a blue plate taken Oct. 1; both plates were taken by C. Brewer and Mueller with the Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. The supernova is located ∼ 3′′ west and 10′′ north of the nucleus of a galaxy at α = 3^{h}59^{m}28^{s}.71, δ = +27^{o}38′47′′.1 (equinox 1950.0). The measured position of the supernova is α = 3^{h}59^{m}28^{s}.47, δ = +27^{o}38′57′′.3. W. Sargent and T. Small obtained spectra using the 5-m Hale Telescope (+ double spectrograph) on Oct. 20 UT, indicating this to be a type-II supernova.

V348 SAGITTARII A. U. Landolt, Louisiana State University; and A. K. Uomoto, The

Johns Hopkins University, observing at Las Campanas Observatory’s 1.0m Swope telescope, report the brightening of the hot R CrB-like star V348 Sgr. V -band CCD images compared with star 12 of Heck et al. (1985, A.Ap. Suppl. 61, 375) yield: Oct. 6.06 UT, V = 16.43; 7.05, 16.26; 8.02, 16.03; 9.06, 15.83; 10.06, 15.58; 11.03, 15.21; 12.02, 14.97; 13.06, 14.61; 14.07, 14.24; 15.07, 13.91; 16.00, 13.62.

COMET HELIN-LAWRENCE (1991l)

Ephemeris extension from orbital elements on MPC 20602:

1992/93 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 25 23^{h}45.^{m}47 +15^{o}55.^{′}3 2.886 3.762 147.^{o}4 8.^{o}2 13.8
Nov. 4 23 34.17 +15 36.0 3.085 3.858 135.8 10.3 14.0
14 23 25.82 +15 19.9 3.309 3.955 124.5 11.9 14.2
24 23 20.18 +15 10.0 3.551 4.051 113.7 12.9 14.4
Dec. 4 23 16.93 +15 08.0 3.806 4.146 103.4 13.4 14.6
14 23 15.68 +15 14.5 4.065 4.241  93.5 13.4 14.8
24 23 16.10 +15 29.6 4.326 4.335  84.0 13.0 15.0
Jan. 3 23 17.87 +15 52.9 4.581 4.429  75.0 12.4 15.2
13 23 20.72 +16 24.0 4.828 4.523  66.2 11.5 15.4
23 23 24.41 +17 02.2 5.063 4.616  57.9 10.4 15.5

1992 October 21 Daniel W. E. Green